# A Comprehensive Survey for Hyperspectral Image Classification: The Evolution from Conventional to Transformers

### v3

@@ 15 Experimental Results and Discussion
Another concern prevalent in recent literature is the overlap between training and test samples [435]. In many cases, training and validation samples are randomly chosen (considering or disregarding the aforementioned point) for the training and validation phases. However, during the testing phase, the entire dataset is often passed through, leading to a highly biased model, as the model has already encountered the training samples, thereby inflating accuracy results. In this study, while the training and test samples are selected randomly (as all models are executed simultaneously), special attention has been given to the above issue. Specifically, measures have been taken to ensure that there is no overlap among these samples, preserving the integrity of the evaluation process. Table 2 furnishes a concise overview of each dataset utilized in the subsequent experiments and further details regarding the experimental datasets are provided in the following.
The University of Houston (UH) dataset, released by the IEEE Geoscience and Remote Sensing Society as part of the Data Fusion Contest in 2013 [434], was collected by the Compact Airborne Spectrographic Imager (CASI). With dimensions of $340\times 1905$ pixels and 144 spectral bands, the dataset features a spatial resolution of $2.5$ meters per pixel (mpp) and a wavelength range from $0.38$ to $1.05$ $\mu$m. The ground truth for this dataset encompasses 15 distinct land-cover classes. Detailed class descriptions and ground truth maps are depicted in Figure 13. Moreover, Table 3 illustrates the selection of class-wise disjoint training, validation, and test samples used to train, validate, and test the models.
The Indian Pines (IP) dataset was acquired by the Airborne Visible/Infrared Imaging Spectrometer (AVIRIS) [432] over the Indian Pines test site in North-western Indiana. Comprising $224$ spectral bands within a wavelength range of $400$ to $2500$ $nm$, the dataset excludes $24$ null and corrupted bands. The image has a spatial size of $145\times{145}$ pixels, with a spatial resolution of 20 mpp. It encompasses $16$ mutually exclusive vegetation classes. Disjoint Training/Test sample maps are illustrated in Figure 14. Moreover, Table 3 reports the selection of class-wise disjoint training, validation, and test samples used to train, validate, and test the models.
@@
The University of Pavia (UP) dataset, captured by the Reflective Optics System Imaging Spectrometer (ROSIS) sensor during a flight campaign over the university campus in Pavia, Northern Italy [433], features dimensions of $610\times{340}$ pixels. With $103$ spectral bands spanning the wavelength range from $430$ to $860~{}{nm}$ and a spatial resolution of 2.5 mpp, the dataset includes 9 urban land-cover classes. Detailed class descriptions and ground truth maps can be found in Figure 15. Moreover, Table 3 illustrates the selection of class-wise disjoint training, validation, and test samples used to train, validate, and test the models.
@@
Table 3: Disjoint Train, Validation, and Test samples selected to train, validate, and test several models.
Indian Pines Dataset University of Houston Dataset Pavia University Dataset
Class Training Validation Test Class Training Validation Test Class Training Validation Test
Alfalfa 4 5 37 Healthy grass 125 125 1001 Asphalt 663 663 5305
Corn-notill 142 143 1143 Stressed grass 125 125 1004 Meadows 1864 1865 14920
Corn-mintill 83 83 664 Synthetic grass 69 70 558 Gravel 209 210 1680
Corn 23 24 190 Trees 124 124 996 Trees 306 306 2452
Grass-pasture 48 48 387 Soil 124 124 994 Painted 134 135 1076
Grass-trees 73 73 584 Water 32 33 260 Soil 502 503 4024
Grass-mowed 2 3 23 Residential 126 127 1015 Bitumen 133 133 1064
Hay-windrowed 47 48 383 Commercial 124 124 996 Bricks 368 368 2946
Oats 2 2 16 Road 125 125 1002 Shadows 94 95 758
Soybean-notill 97 97 778 Highway 122 123 982
Soybean-mintill 245 246 1964 Railway 123 124 988
Soybean-clean 59 59 475 Parking Lot 1 123 123 987
Wheat 20 21 164 Parking Lot 2 46 47 376
Woods 126 127 1012 Tennis Court 42 43 343
Buildings 38 39 309 Running Track 66 66 528
Stone-Steel 9 9 75
To substantiate the outlined concepts in this survey and validate the claims, recent contributions encompass a variety of models, including 2D CNN, 3D CNN, Hyrbid CNN, 2D Inception Network, 3D Inception Net, Hybrid Inception Net, 2D Xception Net, (2+1)D Extreme Exception Net (EX Net), Attention Graph CNN, SCSNet: Sharpend Cosine Similarity-based Neural Network, and Spatial-Spectral Transformer [273]. A comparative analysis of experimental results has been conducted by considering representative works for each model as shown in Table 4, Table 5, Table6 and Figure 16, Figure 17, and Figure 18. To a certain degree, all the aforementioned studies rely on Convolutional Networks and undergo evaluation using three benchmark HSI datasets, namely IP, PU, and UH Scene. This survey specifically focuses on assessing the robustness of these models, taking into account the challenge of classifying HSI with a limited sample size in the training data for joint spatial-spectral classification. Comparative methods tend to misclassify samples with similar spatial structures, exemplified by the confusion between Meadows and Bare Soil classes in the Pavia University dataset, as illustrated in Table 4, Table 5, and Table6 and Figure 16, Figure 17, and Figure 18. Furthermore, the overall accuracy for the Grapes Untrained class is lower compared to other classes, attributed to the reasons mentioned earlier. In summary, it can be concluded that higher accuracy can be attained by augmenting the number of labeled training samples. Therefore, a larger set of labeled training samples has the potential to yield improved accuracies across all competing methods.
In summary, Transformer and Graph-based classification yields significantly better results than convolution-based methods, especially when dealing with the limited availability of labeled training samples [12, 436]. While CNNs can capture the internal structure, the resulting feature representation may lack task-specific characteristics, explaining their comparatively lower performance compared to other models [48, 248, 273, 249].
@@ 16 Conclusions and Future Research Directions
This survey has provided a comprehensive exploration of the evolution from conventional methods to the promising era of Transformers, Explainable AI, and Diffusion models in the domain of Hyperspectral Image Classification (HSC). By thoroughly examining a wide range of methodologies, spanning from TML approaches to state-of-the-art Transformer-based models, our objective was to provide a comprehensive overview of the current landscape and shed light on the potential prospects of HSC. The journey began with an in-depth analysis of conventional methods, including spectral-spatial and statistical approaches. We delved into the strengths and limitations of these conventional models, acknowledging their contributions to the field while recognizing the inherent challenges they face, particularly in handling the complexity of HS data.
As we transitioned into the era of Transformers, we witnessed a paradigm shift in the approach to HSC. The survey meticulously explored the applications of Transformer-based models, drawing attention to the remarkable success achieved by these architectures in various domains. The introduction of self-attention mechanisms, multi-head attention, and positional encoding have empowered Transformers to capture intricate spectral-spatial dependencies, unlocking new possibilities for accurate and context-aware classification.
Our survey also shed light on the challenges that researchers and practitioners encounter in this evolving landscape. From the scarcity of labeled HS data to the interpretability of Transformer-based models, each challenge represents an opportunity for further exploration and innovation. We emphasized the significance of interpretable AI, the role of attention mechanisms, and the ongoing quest for efficient models that can operate in real-time or under resource constraints. In short, this survey endeavors to be a guiding beacon for researchers and practitioners venturing into HSC, providing insights into the past, present, and future of this dynamic field. The fusion of conventional wisdom with transformative technologies promises to elevate the accuracy, efficiency, and interpretability of HSC, paving the way for a new era of discoveries and applications.
The future of HSC is poised for transformative developments, with a focus on advancing Transformer architectures, improving interpretability, and integrating domain knowledge. Anticipated enhancements in HS data acquisition, labeling techniques, and the emergence of real-time, resource-efficient models underscore the evolution of this field. Cross-domain transfer learning and collaborative interdisciplinary research are expected to play pivotal roles, fostering applications in precision agriculture and environmental monitoring. The collective impact of these advancements promises to revolutionize decision-making processes and promote sustainability across diverse domains. The fusion of domain knowledge, advancements in Transformer architectures, and the integration of interpretability measures will shape the next generation of models. As we navigate this transformative journey, collaboration between researchers, domain experts, and technology enthusiasts will be crucial to overcoming challenges and unlocking the full potential of HSI.

### v4

@@ 16 Experimental Results and Discussion
Another concern prevalent in recent literature is the overlap between training and test samples (Liang et al.,, 2017). In many cases, training and validation samples are randomly chosen (considering or disregarding the aforementioned point) for the training and validation phases. However, during the testing phase, the entire dataset is often passed through, leading to a highly biased model, as the model has already encountered the training samples, thereby inflating accuracy results. In this study, while the training and test samples are selected randomly (as all models are executed simultaneously), special attention has been given to the above issue. Specifically, measures have been taken to ensure that there is no overlap among these samples, preserving the integrity of the evaluation process. Table 1 furnishes a concise overview of each dataset utilized in the subsequent experiments and further details regarding the experimental datasets are provided in the following.
The University of Houston (UH) dataset, released by the IEEE Geoscience and Remote Sensing Society as part of the Data Fusion Contest in 2013 (Hong et al., 2021b, ), was collected by the Compact Airborne Spectrographic Imager (CASI). With dimensions of $340\times 1905$ pixels and 144 spectral bands, the dataset features a spatial resolution of $2.5$ meters per pixel (mpp) and a wavelength range from $0.38$ to $1.05$ $\mu$m. The ground truth for this dataset encompasses 15 distinct land-cover classes. Detailed class descriptions and ground truth maps are depicted in Figure 13.
The Indian Pines (IP) dataset was acquired by the Airborne Visible/Infrared Imaging Spectrometer (AVIRIS) (Green et al.,, 1998) over the Indian Pines test site in North-western Indiana. Comprising $224$ spectral bands within a wavelength range of $400$ to $2500$ $nm$, the dataset excludes $24$ null and corrupted bands. The image has a spatial size of $145\times{145}$ pixels, with a spatial resolution of 20 mpp. It encompasses $16$ mutually exclusive vegetation classes. Disjoint Training/Test sample maps are illustrated in Figure 14.
@@
The University of Pavia (UP) dataset, captured by the Reflective Optics System Imaging Spectrometer (ROSIS) sensor during a flight campaign over the university campus in Pavia, Northern Italy Hong et al., 2020b , features dimensions of $610\times{340}$ pixels. With $103$ spectral bands spanning the wavelength range from $430$ to $860~{}{nm}$ and a spatial resolution of 2.5 mpp, the dataset includes 9 urban land-cover classes. Detailed class descriptions and ground truth maps can be found in Figure 15.
@@
To substantiate the outlined concepts in this survey and validate the claims, recent contributions encompass a variety of models, including 2D CNN, 3D CNN, Hyrbid CNN, 2D Inception Network, 3D Inception Net, Hybrid Inception Net, 2D Xception Net, (2+1)D Extreme Exception Net (EX Net), Attention Graph CNN, SCSNet: Sharpend Cosine Similarity-based Neural Network, and Spatial-Spectral Transformer (Ahmad et al., 2024e, ). A comparative analysis of experimental results has been conducted by considering representative works for each model as shown in Table 2, Table 3, Table4 and Figure 16, Figure 17, and Figure 18. To a certain degree, all the aforementioned studies rely on Convolutional Networks and undergo evaluation using three benchmark HSI datasets, namely IP, PU, and UH Scene. This survey specifically focuses on assessing the robustness of these models, taking into account the challenge of classifying HSI with a limited sample size in the training data for joint spatial-spectral classification. Comparative methods tend to misclassify samples with similar spatial structures, exemplified by the confusion between Meadows and Bare Soil classes in the Pavia University dataset, as illustrated in Table 2, Table 3, and Table4 and Figure 16, Figure 17, and Figure 18. Furthermore, the overall accuracy for the Grapes Untrained class is lower compared to other classes, attributed to the reasons mentioned earlier. In summary, it can be concluded that higher accuracy can be attained by augmenting the number of labeled training samples. Therefore, a larger set of labeled training samples has the potential to yield improved accuracies across all competing methods.
In summary, Transformer and Graph-based classification yields significantly better results than convolution-based methods, especially when dealing with the limited availability of labeled training samples (Ahmad et al., 2024d, ; Ahmad et al., 2024g, ). While CNNs can capture the internal structure, the resulting feature representation may lack task-specific characteristics, explaining their comparatively lower performance compared to other models (Ahmad et al., 2021b, ; Ahmad et al., 2024a, ; Ahmad et al., 2024e, ; Butt et al.,, 2024).
@@ 17 Conclusions and Future Research Directions
This survey has charted the evolution of Hyperspectral Image Classification (HSC), covering traditional machine learning (TML), deep learning (DL), the latest Transformer-based architectures, and the Mamba model. Our goal was to provide a well-rounded understanding of the challenges and advancements within the HSC domain. Conventional TML approaches, while foundational, face challenges in handling high-dimensional HS data, prompting a shift towards Transformer and Mamba models with improved spectral-spatial feature handling and interoperability. Through comparative insights and experiments, we highlighted recent progress in Transformer models with attention mechanisms, self-supervised learning, and the use of diffusion models for denoising and feature extraction. The added section on Mamba-based models addresses state-of-the-art approaches, showing potential for integrating spatial-spectral information through architectures tailored to HSC. Key challenges like data scarcity, model interpretability, and real-time performance remain, and future research should focus on domain-specific advancements, enhancing model efficiency, and leveraging cross-domain transfer learning. By providing insights into readily available models, this survey aims to support future research with a comprehensive baseline, encouraging innovations that drive real-world applications in precision agriculture, environmental monitoring, and more.